# Bringing Diversity from Diffusion Models to Semantic-Guided Face Asset Generation

### v1

@@ Abstract.
Digital modeling and reconstruction of human faces serve various applications. However, its availability is often hindered by the requirements of data capturing devices, manual labor, and suitable actors. This situation restricts the diversity, expressiveness, and control over the resulting models. This work aims to demonstrate that a semantically controllable generative network can provide enhanced control over the digital face modeling process. To enhance diversity beyond the limited human faces scanned in a controlled setting, we introduce a novel data generation pipeline that creates a high-quality 3D face database using a pre-trained diffusion model. Our proposed normalization module converts synthesized data from the diffusion model into high-quality scanned data. Using the 44,000 face models we obtained, we further developed an efficient GAN-based generator. This generator accepts semantic attributes as input, and generates geometry and albedo. It also allows continuous post-editing of attributes in the latent space. Our asset refinement component subsequently creates physically-based facial assets. We introduce a comprehensive system designed for creating and editing high-quality face assets. Our proposed model has undergone extensive experiment, comparison and evaluation. We also integrate everything into a web-based interactive tool. We aim to make this tool publicly available with the release of the paper.
@@ 1. Introduction
Creating realistic 3D human faces has been a long-standing goal in both the digital industry and academia. Differently from a 2D portrait, 3D face assets contain detailed information on skin texture, facial geometry, and materials in order to be applicable in the physically based rendering (PBR) pipeline. The demands for them are rapidly growing in many of today’s digital industries - gaming, movie, teleconference, and AR/VR, to list a few. However, traditionally, creating such an asset in modern industrial practices requires a time-consuming process from professional artists and, in the case of 3D scans, a meticulously calibrated facial capture studio. As of today, high-quality 3D face assets are still very costly to create.
Naturally, many past academia and industry, reasearch and engineering efforts have been made to automatize the creation of high quality 3D human faces, in a hope to make this technology more accessible to small studios and individuals. The scarcity of available face asset dataset, however, posts major challenges to even the current state-of-the-art face generation models. Industrial toolkits such as the MetaHuman (Fang et al., 2021), for example, are based on preset face assets that restrict the modeling capabilities to discrete data points and necessitates mannual editing with professional artist knowledge. Conventional 3D Morphable Models (3DMMs) (Paysan et al., 2009; Booth et al., 2018; Cao et al., 2014; Li et al., 2017) and learning-based methods (Li et al., 2020b; Yang et al., 2020; Zhang et al., 2023a) have two main limitations: they struggle to effectively integrate textures with underlying geometry, and their parametric spaces fail to capture the full spectrum of human facial geometry (Liu et al., 2022). Additionally, these models lack robust semantic and attribute controls for generation.
The issue of data scarcity becomes more pronounced when considering modeling the diverse distribution of not just facial geometries, but also facial textures that encode variations of skin tone, freckles, wrinkles and even pores, etc. In DreamFace (Zhang et al., 2023a), a highly relevant work to ours, the proposed model has been shown to create 3D face assets with high quality, but we observe that the diversity of the synthesized textures are limited and biased to Asian skin due to the database bias (over 90% of FaceScape (Yang et al., 2020) is Asian), as only 618 textures are available for them to fine-tune a diffusion model, which we believe is insufficient. While editing features like tattoos and makeup can be achieved using a pre-trained texture LDM with prompts, this approach only allows for single-pass editing without precise control and does not support continuous modifications. Increasing texture diversity of synthesized digital humans is a key motivation under many recent 3D generative models that leverage the vision-language model CLIP (Radford et al., 2021) (as seen in AvatarClip (Hong et al., 2022)) or large diffusion model (as seen in DreamAvatar (Cao et al., 2023)) as guidance. As CLIP and many large diffusion models learn from in-the-wild data that is larger in size by some orders of magnitude, knowledge distillation from them towards digital human generation has been shown effective, particularly in the much increased diversity of the generated textures. In the more general domain, LucidDreamer (Liang et al., 2023) addresses this issue by introducing Interval Score Matching (ISM), which produces high-quality 3D models with more details and better multi-view consistency. However, 3D face generated from the above methods often contain artifacts and incomplete assets, and are not in a format readily usable for animation and rendering. Additionally, the diffusion process remains slow.
Our efforts in face asset generation similarly aim to address the data limitation challenge, by leveraging the expressive power of a large diffusion model to diversely synthesize these assets. However, we have added three key design layers that distinguish our framework from previous 3D face generation models and bridge the quality gap between scanned high-quality data and mass-generated synthetic data: 1) UV textures estimated from diffusion-generated portraits is by itself incomplete and of arbitrary shading, which are adapted poorly to the PBR pipeline. We devise a novel method to process these reconstructed UV textures through texture completion and a normalization step that produces clean, complete albedo maps from the raw textures. 2) We additionally take careful considerations in supporting meaningful semantic control and editing of the generated 3D faces. While the semantic labels are extracted from a pretrained text-guided diffusion model, we propose several designs that greatly improve our final model’s ability to both control and edit its generated faces based on a certain facial attributes, when compared to the previous works that also incorporated this feature (Zhou et al., 2024; Wu et al., 2023; Zhang et al., 2023a). The first step taken is to formulate the semantic attributes with the demographic properties: ethnicity, gender, and age. This circumvents the ambiguity of using a general text prompt and language model to interpret user inputs and provides more accurate control. The second step replaces the common practices of conditioning the attributes in generation with a disentangled framework based on the generative adversarial network (GAN) (Goodfellow et al., 2014a). In this framework, a two-step adversarial training scheme, extended from (Xiang et al., 2021), is designed to learn an encoder that extracts unlabeled information (e.g. identity), and a generator that produces semantically accurate image from labels and the unlabeled information. This design not only circumvents the difficulty of inversion with a diffusion model, but also is validated by experiments (see Figure 13 and results in supplementary video) to be more effective in preserving the identity of the face image that is being edited. 3) Last but not least, efficiency is a key factor in measuring the accessibility of a face asset generation model. While traditional approaches that scans, reconstructs and cleans up a captured face may take days to produce a complete face asset (LeGendre et al., 2018), recent AI-based methods (Liu et al., 2022; Zhang et al., 2023a; Zhou et al., 2024) have greatly accelerated this process. Controlled 3D face generation by sampling from a latent diffusion model is typically slow due to its iterative nature. By distilling information to a GAN network, our generation model achieves a drastic speed up in generating shape and albedo (0.014s vs. 40s under a single Nvidia A6000 GPU).
@@
Our training pipeline consists of two major stages. The first stage leverages a image-based diffusion network and a 3D face reconstruction network to create a dataset of diverse, high-quality 3D face assets. As the direct image output of the diffusion network contains lighting and other external visual effects, albedo maps are computed from a proposed texture normalization algorithm. In addition, we apply a sanity check to ensure quality and discard any data with artifacts or mismatched labels, and convert the conditioning text used to generate the data into three controlling attributes: age, gender, and ethnicity. This process resulted in a dataset of 44k set of 3D face models. The second stage extends from DisUnknown (Xiang et al., 2021) to train a GAN with the processed training data from the first stage. As described, this stage produces the final model that generates 3D face assets consisting of 4k geometry, albedo, specular and displacement maps in the UV space. Thanks to training under a disentanglement objective, the flexible model can either create 3D assets from an attribute description (“generate a 40 years-old, Hispanic male”), or perform inversion on a given image and subsequently edit and reconstruct a 3D asset from the image, based on an attribute description (“turn this face photo into a 3D face of a 40 years-old, Hispanic male, without losing their identity”).
We summarize our contributions as follow:
@@ 2.2. Text-to-3D Face
Advancements in text-to-3D avatar generation (Michel et al., 2022; Hong et al., 2022; Wu et al., 2023) have leveraged the vision-language model CLIP (Radford et al., 2021) to map natural language descriptions into latent spaces, enabling the generation of 2D images that are then converted into 3D avatars. While this approach allows attribute and semantic control, it often suffers from texture artifacts and inconsistent multi-view outputs. Beyond avatar generation, text-to-3D content creation in general has rapidly evolved, with recent works (Cao et al., 2023; Chen et al., 2023a; Lin et al., 2023; Metzer et al., 2023; Poole et al., 2022) excelling at generating diverse 3D models from text. These methods employ pre-trained text-to-image diffusion models as priors to guide the training of parameterized 3D models, ensuring multi-view consistency and text alignment through Score Distillation Sampling (SDS). However, SDS encourages average-seeking behaviors in the 3D representation, leading to over-smoothed results. LucidDreamer (Liang et al., 2023) addresses this issue with Interval Score Matching (ISM), improving detail quality, but their method remains computationally expensive — taking approximately 36 minutes to generate a 3D model.
@@
Inspired by DreamFusion (Poole et al., 2022) and LucidDreamer (Liang et al., 2023), we aim to refine generative 3D generation. While these models generate 3D faces from text, they often produce artifacts and require manual quality control, making them unsuitable for direct application. Nonetheless, pre-trained image diffusion models, trained on large datasets, naturally enhance diversity, provide annotations and offer great accessibility. Our approach harnesses the rich priors of image diffusion models to improve quality, control, and usability in 3D avatar generation.
@@ 2.3. Semantic Face Generation and Manipulation
Generative Adversarial Networks (GANs), pioneered by (Goodfellow et al., 2014b), have been well studied in the setting of semantic editing. A key challenge lies in achieving semantic and disentangled control in generative models, e.g., randomly changing one specific attribute while preserving the other attributes. Efforts to interpolate in the latent space for smooth output variation have been explored by (Laine, 2018; Shao et al., 2018), while recent research focuses on disentangling the latent space for semantic control (Shen et al., 2020; Härkönen et al., 2020; Shen and Zhou, 2020; Jiang et al., 2021; Zheng et al., 2021). Observations of vector arithmetic in latent space by (Radford et al., 2016; Upchurch et al., 2017) have led to unsupervised disentanglement methods. Semi-supervised methods apply principal component analysis for attribute identification (Härkönen et al., 2020), while supervised methods use labeled data for latent space factorization (Shen et al., 2020; Kowalski et al., 2020). Alternative approaches utilize 3DMM for semantic control over pre-trained StyleGAN latent space (Tewari et al., 2020a; Wang et al., 2022) or train unsupervised data with labels from attribute classifiers (Khodadadeh et al., 2022). Also, to transfer such controllability to real image editing, the GAN inversion methods (Abdal et al., 2019; Zhu et al., 2020; Tewari et al., 2020b) propose to map the real image into the latent space and thus can manipulate real images.
@@ 3. Method
Our goal is to create a 3D avatar synthesis model that uses attribute prompts (ethnicity, gender and age) to produce high-quality 3D face assets, including 3D geometry in the form of a position map and detailed textures in the forms of albedo, specular and displacement maps. To achieve our goal, we start by constructing a large, high-quality 3D face dataset using a pre-trained diffusion model. We introduce a unique framework in Section 3.1 to ensure the collected data are clean and complete with corresponding labels. Then, a dedicated generative network, as detailed in Section 3.2, is trained using these pairs of facial data to create the basic geometries and albedo textures. Lastly, we apply post-processing, as described in Section 3.3, to refine these basic assets and generate PBR assets.
@@ 3.1. Data Preparation
We leverage the rich priors of a large diffusion model by using it to synthesize a dataset of attribute-controlled faces. However, existing large diffusion model is limited to generating 2D portraits, does not provide identity control, and the generated portraits may not conform to the provided labels. This section details four stages needed to create 3D assets (shape and albedo texture) from the raw diffusion-generated 2D portraits. First, we use the pre-trained latent diffusion models (Rombach et al., 2022; Zhang et al., 2023b) to create a set of portraits based on semantic text and geometric information. Next, these portraits are projected onto the UV space associated with the topology of the underlying 3D surface, yielding an incomplete UV-space texture, and the geometries are refined with a single-image 3D reconstruction network. In the third stage, the incomplete UV-space textures are refined using a texture completion algorithm, and the underlying geometry is optimized to ensure consistency with the synthesized facial appearances. Lastly, we introduce a novel normalization network, trained on data from two domains - a high-quality scanned database and the generated textures dataset. It converts unconstrained textures into clean albedo, similar to the scanned data domain. This pipeline efficiently creates a high-quality, diverse, and realistic 3D facial avatar dataset, including subjects, suitable for subsequent generator training. The data generation pipeline is illustrated in Figure 2.
@@ 3.1.1. Portraits Generation using Diffusion Models
We utilize pre-trained checkpoints from Stable Diffusion v1.5 (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) (pretrained on normal maps) to generate portraits with precise control over semantic facial attributes and geometric information. To avoid ambiguity caused by the language model and ensure precise control, we define three major demographic attributes that contribute most to human appearances and shapes: ethnicity, gender, and age. We also include generic prompts for illumination and quality control, like resolution and framing. For example, ”East Asian female age 20” is one set of demographic attribute combinations. Age groups span from 15 to 75 years old. For gender, we use the labels male, female, and unisex. We also consider ethnicity/race and geographic location, resulting in 14 distinct categories: East Asian, Southeast Asian, Middle Eastern, Caucasian, Germanic, Celtic, Slavic, Romance, Australian, Native Americans, Aboriginal, Pacific Islander, and African. To provide guidance to the generation, the pretrained ControlNet is used to condition the diffusion process on a normal map, rendered at the front view from a 3D face randomly sampled from a scanned dataset. 3D geometry is obtained by using a single-view face reconstruction model, a variant of ReFA (Liu et al., 2022), on the 2D portraits. In total, 65k portraits with 3D geometry are created from these steps. Ethnicity is uniformly sampled from 14 categories. As for gender, the distribution consists of male, female, and unisex. Age is uniformly sampled from the following groups: , , , , , , , , , , , , and .
@@ 3.1.2. Texture Completion
With the corresponding geometries, the 2D portraits can be directly resampled to the corresponding UV space. Figure 2 illustrates the projected portraits. However, since only the frontal part of the face is visible, the missing regions in the boundary of the UV maps need to be completed. We first search for the nearest neighbor with the highest similarity to the projected portrait in a synthesized texture database, using Peak Signal-to-Noise Ratio (PSNR) as our measure. This database is derived from our scanned database, and we use it to find a reference to fill missing regions. It consists of UV-space images rendered using geometry and textures from the scanned database, combined with randomized lighting, as demonstrated in the PBR texture maps shown in Figure 2. The complete UV-space texture is then obtained by blending the projected portraits and texture maps with the pyramid blending algorithm (Burt and Adelson, 1983). Finally, a sanity check is applied to filter out results with artifacts to maintain high data quality. The filtering step keeps approximately UV textures from the original portraits. However, these textures still contain baked-in lighting, which are not part of our desired clean albedo. Therefore, a further normalization step is proposed in the next section.

### v2

@@ Abstract.
High-quality 3D face asset creation remains costly due to reliance on controlled capture setups and manual processing, limiting scalability and diversity. We introduce a fully automated, semantically controllable framework for generating PBR-ready 3D facial assets without requiring dedicated scans. Our pipeline begins with a diffusion-based data synthesis stage, where 2D portrait samples from a pre-trained diffusion model are converted into 44K textured 3D face reconstructions via our proposed geometry recovery and texture normalization algorithm, which aligns arbitrarily shaded outputs into clean albedo space. Using this dataset, we train a disentangled adversarial generator that maps semantic attributes (age, gender, ethnicity) to UV-space geometry and albedo, enabling both direct sampling and continuous latent editing while preserving identity. A refinement stage further produces PBR materials and secondary assets (eyeballs, teeth, gums). The resulting system supports controllable face generation and post-editing in real time and exports directly to standard rendering and animation pipelines. We evaluate each component extensively and provide a web-based interactive interface to showcase practical deployment.
@@ 1. Introduction
Creating realistic 3D human faces has long been a central goal in both industry and academia. A high-quality 3D face asset integrates detailed skin textures, accurate geometry, and material properties to function within a physically based rendering (PBR) pipeline. Demand for such assets continues to surge across gaming, film, teleconferencing, and AR/VR applications. However, producing them remains labor-intensive and expensive—typically requiring skilled artists and, in the case of scans, access to a carefully calibrated facial capture studio. Even today, generating production-ready 3D face assets is still a costly and time-consuming endeavor.
Despite extensive efforts to automate high-quality 3D face creation for broader accessibility, current face generation models remain constrained by the scarcity and imbalance of available facial asset datasets. Industrial systems like MetaHuman (Fang et al., 2021) rely on limited preset assets, while traditional 3D Morphable Models (Paysan et al., 2009; Booth et al., 2018; Cao et al., 2014; Li et al., 2017) and learning-based approaches (Li et al., 2020b; Yang et al., 2020; Zhang et al., 2023a) struggle to jointly model geometry and texture, failing to capture the full diversity of real human faces or offer precise semantic control. Data limitations are especially evident in facial texture generation—skin tone, freckles, wrinkles, and pores are poorly represented, leading to biased outputs such as DreamFace (Zhang et al., 2023a), which predominantly synthesizes Asian-like textures due to dataset skew (over 90% of FaceScape (Yang et al., 2020) is Asian). Recent methods that leverage CLIP (Radford et al., 2021) or large diffusion models improve diversity through knowledge distillation, as seen in AvatarClip (Hong et al., 2022) and DreamAvatar (Cao et al., 2023), while LucidDreamer (Liang et al., 2023) enhances detail using Interval Score Matching. However, these approaches still suffer from artifacts, incomplete assets, slow inference, and lack of animation-ready outputs, leaving a significant gap between generative quality and production usability.
Our approach addresses data scarcity in 3D face asset generation by leveraging large diffusion models for diverse synthesis while introducing three key innovations that bridge the gap between scanned-quality assets and synthetic outputs. (1) Diffusion-generated UV textures are often incomplete and inconsistently shaded, making them unsuitable for PBR rendering. We introduce a texture completion and normalization pipeline to produce clean, fully albedo-compatible maps. (2) We enable precise semantic control and editing by structuring attributes around demographic labels (ethnicity, gender, age) and adopting a disentangled GAN-based framework, which preserves identity during editing more effectively than diffusion-based inversion. (3) We drastically reduce generation time by distilling knowledge into a GAN, achieving 0.014s versus 40s per face compared to a diffusion model on a single A6000 GPU—making high-quality face asset generation both controllable and real-time.
@@
Our training pipeline consists of two major stages. The first stage leverages an image-based diffusion network and a 3D face reconstruction network to create a dataset of diverse, high-quality 3D face assets. As the direct image output of the diffusion network contains lighting and other external visual effects, albedo maps are computed from a proposed texture normalization algorithm. In addition, we apply a sanity check to ensure quality and discard any data with artifacts or mismatched labels, and convert the conditioning text used to generate the data into three controlling attributes: age, gender, and ethnicity. This process resulted in a dataset of 44K 3D face models. The second stage extends from DisUnknown (Xiang et al., 2021) to train a GAN with the processed training data from the first stage. As described, this stage produces the final model that generates 3D face assets consisting of 4K geometry, albedo, specular and displacement maps in the UV space. Thanks to our unique design on adversarial training, the flexible model can either create 3D assets from an attribute description (“generate a 40-year-old, Hispanic male”), or perform inversion on a given image and subsequently edit and reconstruct a 3D asset from the image, based on an attribute description (“turn a face photo into a 3D face of a 40-year-old, Hispanic male, without losing their identity”).
We summarize our contributions as follows:
@@ 2.2. Text-to-3D Face
Advancements in text-to-3D avatar generation (Michel et al., 2022; Hong et al., 2022; Wu et al., 2023) have leveraged the vision-language model CLIP (Radford et al., 2021) to map natural language descriptions into latent spaces, enabling the generation of 2D images that are then converted into 3D avatars. While this approach allows attribute and semantic control, it often suffers from texture artifacts and inconsistent multi-view outputs. Beyond avatar generation, text-to-3D content creation in general has rapidly evolved, with recent works (Cao et al., 2023; Chen et al., 2023a; Lin et al., 2023; Metzer et al., 2023; Poole et al., 2022) excelling at generating diverse 3D models from text. These methods employ pre-trained text-to-image diffusion models as priors to guide the training of parameterized 3D models, ensuring multi-view consistency and text alignment through Score Distillation Sampling (SDS). However, SDS encourages average-seeking behaviors in the 3D representation, leading to over-smoothed results. LucidDreamer (Liang et al., 2023) addresses this issue with Interval Score Matching (ISM), improving detail quality, but their method remains computationally expensive — taking approximately 36 minutes to generate a 3D model. A challenge in text-to-3D generation is that the appearance often entangles reflectance and illumination. Intrinsic decomposition addresses this issue by separating reflectance (albedo) from lighting, and diffusion priors (Chen et al., 2024; Luo et al., 2024; Kocsis et al., 2024; Li et al., 2025) have recently been used to improve such decomposition in a general image. At the same time, text-to-3D face avatar generation requires producing a usable 3D facial asset, which imposes additional constraints than those for general images.
@@
Inspired by DreamFusion (Poole et al., 2022), LucidDreamer (Liang et al., 2023) and HeadStudio (Zhou et al., 2024b), we aim to refine generative 3D generation. While these models generate 3D faces from text, they often produce artifacts and require manual quality control, making them unsuitable for direct application. Nonetheless, pre-trained image diffusion models, trained on large datasets, naturally enhance diversity, provide annotations and offer great accessibility. Our approach harnesses the rich priors of image diffusion models to improve quality, control, and usability in 3D avatar generation.
@@ 2.3. Semantic Face Generation and Manipulation
Generative Adversarial Networks (GANs), pioneered by (Goodfellow et al., 2014), have been well studied in the setting of semantic editing. A key challenge lies in achieving semantic and disentangled control in generative models, e.g., randomly changing one specific attribute while preserving the other attributes. Efforts to interpolate in the latent space for smooth output variation have been explored by (Laine, 2018; Shao et al., 2018), while recent research focuses on disentangling the latent space for semantic control (Shen et al., 2020; Härkönen et al., 2020; Shen and Zhou, 2020; Jiang et al., 2021; Zheng et al., 2021). Observations of vector arithmetic in latent space by (Radford et al., 2016; Upchurch et al., 2017) have led to unsupervised disentanglement methods. Semi-supervised methods apply principal component analysis for attribute identification (Härkönen et al., 2020), while supervised methods use labeled data for latent space factorization (Shen et al., 2020; Kowalski et al., 2020). Alternative approaches utilize 3DMM for semantic control over pre-trained StyleGAN latent space (Tewari et al., 2020a; Wang et al., 2022) or train unsupervised data with labels from attribute classifiers (Khodadadeh et al., 2022). Also, to transfer such controllability to real image editing, the GAN inversion methods (Abdal et al., 2019; Zhu et al., 2020; Tewari et al., 2020b) propose to map the real image into the latent space and thus can manipulate real images. Recently, (Fadaeinejad et al., 2025) incorporated artist input into a GAN-based framework, enabling greater controllability in the avatar creation workflow.
@@ 3. Method
Our goal is to create a 3D avatar synthesis model that uses facial attributes (ethnicity, gender and age) to produce high-quality 3D face assets, including 3D geometry and PBR-based material properties such as albedo, specular and displacement maps. To achieve our goal, we start by constructing a large, high-quality and diverse 3D face assets dataset associated with corresponding annotations using a pre-trained diffusion model. We introduce a unique framework in Section 3.1 to construct a 3D face asset dataset that is clean and complete with corresponding labels. Then, a dedicated generative network, as detailed in Section 3.2, is trained using the obtained dataset to create the basic geometries and albedo textures. Lastly, we apply post-processing, as described in Section 4.1, to refine the generated face geometries and albedo textures and generate the specular and displacement maps to complete the PBR assets.
@@ 3.1. Data Preparation
We leverage the rich priors of an image diffusion model by using it to synthesize a dataset of high-quality 3D faces with annotations, which is used to train our Text2Avatar 3D face asset generator. This section details three main steps needed to create a dataset of attribute-controlled 3D assets (shapes and albedo textures). Section 3.1.1 describes the generation of 2D portraits that leverages information from both the rich priors of a large image diffusion model (Rombach et al., 2022; Zhang et al., 2023b) and a dataset of scanned 3D faces from LightStage (Li et al., 2020b) and the TrippleGangers dataset (Triplegangers, 2021). Section 3.1.2 describes the construction of 3D face geometries and complete UV textures from the generated portraits. Section 3.1.3 details a normalization method that removes lighting and other external visual effects from the textures to create the clean albedo maps.
@@ 3.1.1. Portraits Generation using Diffusion Models
We utilize the pre-trained text-to-image Stable Diffusion v1.5 (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) (trained on normal maps) to generate a rich collection of 2D portraits. Specifically, the diffusion model is given various text descriptions of facial attributes. To avoid ambiguity caused by the language model and ensure precise control, we define three major demographic attributes that contribute most to human appearances and shapes: ethnicity, gender, and age. We also include generic prompts for illumination and quality control, like resolution and framing. For example, ”East Asian female age 20” is one set of demographic attribute combinations. Age groups span from 15 to 75 years old. For gender, we use the labels male, female, and unisex (a gender-neutral semantic category positioned between male and female). We also consider ethnicity/race and geographic location, resulting in 14 distinct categories: African, East Asian, Southeast Asian, South Asian, Middle Eastern, Caucasian, Germanic, Celtic, Slavic, Romance, Australian, Native Americans, Aboriginal, and Pacific Islander. In addition, we condition the portrait generation with normal maps, rendered from a randomly sampled face geometry at the frontal view, from the scan dataset . The normal maps provide head pose and geometry guidance to constrain the synthesized image.
Since random sampling may create a conflict between the desired attributes and the selected geometry, we further reconstruct a refined face geometry based on the alignment between the generated portrait and the sampled face geometry using ReFA (Liu et al., 2022) (see Section 4.1 for details).
In total, 65K portraits with 3D geometry are created at this stage. Ethnicity is uniformly sampled from 14 categories. As for gender, the distribution consists of male, female, and unisex. Age is uniformly sampled from the following groups: , , , , , , , , , , , , and .
@@ 3.1.2. Texture Completion
The first step produces a collection of 2D portraits, each with the corresponding attribute labels and predicted geometry. To process this data into 3D assets, we complete the UV-space texture maps from these outputs. Specifically, the texture map is initialized by directly projecting the 2D portraits based on the predicted geometry. However, since only the frontal part of the face is visible, we propose a completion method to fill in the missing regions (e.g. ears) in the boundary of the UV maps. We first collect a dataset of skin textures by baking randomized lighting into the UV-space textures of the scan dataset (see Figure 2). Then, given the projected partial textures, we search for the nearest neighbor with the highest similarity to the projected portrait in , using Peak Signal-to-Noise Ratio (PSNR) as our measure. The complete UV-space texture is then obtained by blending the projected partial textures and the retrieved texture maps with the pyramid blending algorithm (Burt and Adelson, 1983).
Finally, to ensure the quality of the completed textures, a sanity check is applied to filter out results with artifacts to maintain high data quality. The filtering step keeps approximately UV textures from the original portraits. However, these textures still contain baked-in lighting, which is not part of our desired clean albedo. Therefore, a further normalization step is proposed in the next section.